# Unsupervised Corpus Poisoning Attacks in Continuous Space for Dense Retrieval

### v1

@@ 1. Introduction
Moreover, most attacks in dense retrieval commonly make the assumption of a target query at the time of attack, that is used to inform how documents are corrupted. To name a few, AGGD(Su et al., 2024), IDEM(Chen et al., 2023), PRADA (Wu et al., 2023), PAT (Liu et al., 2022), MCARA (Liu et al., 2023b), TARA (Liu et al., 2023a), as well as the aforementioned corpus poisoning attacks. We claim that this is a strong assumption from a practical standpoint, as we cannot always rely on knowing the target queries in advance, as well as from a scientific standpoint, as it raises concerns of overfitting the attack models on specific queries. In this paper, we introduce a more realistic and challenging scenario, termed unsupervised corpus poisoning (Figure 1). In this context, “unsupervised” refers to the fact that there is no prior knowledge about the query distribution at the time of attack, and the attack method itself is only informed by the target document.
@@ 2. Related work
Word Substitution Attacks: Our method belongs to the family of word substitution ranking attacks (Wu et al., 2022), similar to PRADA (Wu et al., 2023), MCARA (Liu et al., 2023b) or TARA (Liu et al., 2023a), as well as the family of corpus poisoning attacks such as Order-Disorder (Liu et al., 2022) and HotFlip (Ebrahimi et al., 2018; Zhong et al., 2023; Su et al., 2024; Li et al., 2025). However, these methods assume prior knowledge about a query distribution that guides training, and perturbation search. Our work introduces the novel, unsupervised setting, and is uniquely positioned separately from these methods, as we make no assumptions about what kind of queries correspond to the attacked rankings/corpus. The only training signal for our attack method is the document itself. To contextualize our work within the existing literature, we mention that our work is in a similar direction as Zhong et al. (2023), except (i) no queries are used during training, (ii) perturbation search happens in the embedding space, and (iii) our approach is significantly faster and generates adversarial documents with lower perplexity.
@@
Imperceptibility of Attacks: There is a line of adversarial IR work that corrupts documents with word substitution attacks (Wu et al., 2023; Liu et al., 2023b, 2024, a), aiming to maintain imperceptibility. This imperceptibility refers to preserving the original document’s semantics while boosting its ranking during attacks. However, in agreement with (Chen et al., 2022), we argue that imperceptibility is neither necessary nor practical from the user’s perspective. This is because an attack succeeds only when the user actively reads the top-ranked adversarial document and perceives it as nonsensical, We found this assumption to be compatible with the “realistic attack” angle of our work.
@@ 3.3. Adversarial Generation Optimization
where $P\left(\tilde{d}\mid e_{d};\varphi\right)\in\mathbb{R}^{\left|d\right|\times\left|V\right|}$ is predicted probabilities of each token in the adversarial document $\tilde{d}$, abbreviated as $P\left(\tilde{d}\right)$. The perturbation model $\varphi(\cdot)$ is optimized to minimize the following loss function:
@@
In this paper, we use a three-layer perceptron to implement the perturbation model. Notably, although the perturbation model described above is document-specific, meaning a unique attack model is initialized for generating each adversarial document, its lightweight design with few parameters allows for fast and efficient training. Throughout the optimization process using the $L_{Attack}$ loss, the model updates prediction probabilities $P\left(\tilde{d}\right)$ continuously and we can select tokens of the output adversarial document by $\tilde{d}=\operatorname*{arg\,max}_{p(\tilde{t}_{i})\forall\tilde{t}_{i}\in V}P\left(\tilde{d}\right)$.
@@
TREC DL 19 Passage-Retrieval Misc. × Binary 8,841,823 532,761 43 95.4 5.96 55.98
TREC DL 20 Passage-Retrieval Misc. × Binary 8,841,823 532,761 54 66.8 5.96 55.98
@@ 4.1.1. Datasets
For adversarial attacks, we select TREC DL 19 (Craswell et al., 2020b) and TREC DL 20 (Craswell et al., 2020a) as in-domain target datasets as they share the same corpus with MS MARCO. Additionally, we select NQ, Quora (Iyer et al., 2017), FiQA (Maia et al., 2018), and Touché-2020 (Bondarenko et al., 2020) from the BEIR (Thakur et al., 2021) benchmark as out-of-domain datasets due to the diverse performance of retrieval models across these datasets. The statistics of these datasets are shown in Table 1.
@@ 4.1.3. Baseline Methods
HotFlip-based: It is one of the most widely used gradient-based word substitution methods and serves as a foundational technique for many other approaches in generating adversarial documents (Zhong et al., 2023; Wu et al., 2023; Liu et al., 2023b; Su et al., 2024; Liu et al., 2023a). In this paper, we refer to Zhong et al. (Zhong et al., 2023) and use the codebase from Li et al. (Li et al., 2025), as it is an accelerated version.
@@ 4.1.5. Implementation Details
To train the reconstruction model, we use a learning rate of 1e-5, train for 5 epochs, and set the batch size to 128$\times$4 GPUs, with a maximum text length of 128. For all attack experiments, we set $\lambda=5$, the learning rate is 5e-4. We run both our method and the HotFlip-based methods for 3000 epochs. The HotFlip-based adversarial document is initialized as a [MASK] token list of the same length as the target document. Our experiments are mainly implemented using Pytorch 2.1 on a Ubuntu server with Nvidia L4 ×24G × 8 GPUs, AMD EPYC 9554P CPU, and 384G memory. Our code is available at https://github.com/liyongkang123/unsupervised_corpus_poisoning.
@@
Retrievers Accuracy\,$\uparrow$ Precision\,$\uparrow$ Recall\,$\uparrow$ F1\,$\uparrow$
@@
Datasets Attack Methods Embedding Similarly Sematic Irrelevance Perplexity\,$\downarrow$ Time Cost (s)\,$\downarrow$
ASR\,$\uparrow$ Top@10\,$\uparrow$ Top@50\,$\uparrow$ BLEU\,$\downarrow$ LLM Q1\,$\uparrow$ LLM Q2\,$\uparrow$
TREC DL 19 nDCG@10$=$0.650 Random Noise 0.735 0.023 0.046 0.037 1.000 0.984 3469.2 0.1
@@ 4.3.2. Black-box Attack Performance
It is worth noting that, in the literature (Wu et al., 2023; Liu et al., 2023a, 2024, b), black-box attacks typically involve distilling a surrogate model first, then using white-box methods to attack that surrogate model. However, distilling LLMs like LLaMA is too computationally intensive in our experiments. Therefore, to ensure fairness across all target models and to reduce the computational workload, we refrain from training surrogate models. Instead, we follow the approach in  (Li et al., 2025) and directly use the adversarial documents generated by various attack methods during attacks on SimLM and apply them to the target models.

### v2

@@ 1. Introduction
Moreover, most attacks in dense retrieval commonly make the assumption of a target query at the time of attack, that is used to inform how documents are corrupted. To name a few, AGGD(Su et al., 2024), IDEM(Chen et al., 2023), PRADA (Wu et al., 2023), PAT (Liu et al., 2022), MCARA (Liu et al., 2023a), TARA (Liu et al., 2023b), as well as the aforementioned corpus poisoning attacks. We claim that this is a strong assumption from a practical standpoint, as we cannot always rely on knowing the target queries in advance, as well as from a scientific standpoint, as it raises concerns of overfitting the attack models on specific queries. In this paper, we introduce a more realistic and challenging scenario, termed unsupervised corpus poisoning (Figure 1). In this context, “unsupervised” refers to the fact that there is no prior knowledge about the query distribution at the time of attack, and the attack method itself is only informed by the target document.
@@ 2. Related Work
Word Substitution Attacks: Our method belongs to the family of word substitution ranking attacks (Wu et al., 2022), similar to PRADA (Wu et al., 2023), MCARA (Liu et al., 2023a) or TARA (Liu et al., 2023b), as well as the family of corpus poisoning attacks such as Order-Disorder (Liu et al., 2022) and HotFlip (Ebrahimi et al., 2018; Zhong et al., 2023; Su et al., 2024; Li et al., 2025). However, these methods assume prior knowledge about a query distribution that guides training, and perturbation search. Our work introduces the novel, unsupervised setting, and is uniquely positioned separately from these methods, as we make no assumptions about what kind of queries correspond to the attacked rankings/corpus. The only training signal for our attack method is the document itself. To contextualize our work within the existing literature, we mention that our work is in a similar direction as Zhong et al. (2023), except (i) no queries are used during training, (ii) perturbation search happens in the embedding space, and (iii) our approach is significantly faster and generates adversarial documents with lower perplexity.
@@
Imperceptibility of Attacks: There is a line of adversarial IR work that corrupts documents with word substitution attacks (Wu et al., 2023; Liu et al., 2023a, 2024, b), aiming to maintain imperceptibility. This imperceptibility refers to preserving the original document’s semantics while boosting its ranking during attacks. However, in agreement with (Chen et al., 2022), we argue that imperceptibility is neither necessary nor practical from the user’s perspective. This is because an attack succeeds only when the user actively reads the top-ranked adversarial document and perceives it as nonsensical, We found this assumption to be compatible with the “realistic attack” angle of our work.
@@ 3.3. Adversarial Generation Optimization
[…]
@@
[…]
@@
TREC DL 19 Passage-Retrieval Misc. × 4-level 8,841,823 532,761 43 95.4 5.96 55.98
TREC DL 20 Passage-Retrieval Misc. × 4-level 8,841,823 532,761 54 66.8 5.96 55.98
@@ 4.1.1. Datasets
For adversarial attacks, we select TREC DL 19 (Craswell et al., 2020a) and TREC DL 20 (Craswell et al., 2020b) as in-domain target datasets as they share the same corpus with MS MARCO. Additionally, we select NQ, Quora (Iyer et al., 2017), FiQA (Maia et al., 2018), and Touché-2020 (Bondarenko et al., 2020) from the BEIR (Thakur et al., 2021) benchmark as out-of-domain datasets due to the diverse performance of retrieval models across these datasets. The statistics of these datasets are shown in Table 1.
@@ 4.1.3. Baseline Methods
HotFlip-based: It is one of the most widely used gradient-based word substitution methods and serves as a foundational technique for many other approaches in generating adversarial documents (Zhong et al., 2023; Wu et al., 2023; Liu et al., 2023a; Su et al., 2024; Liu et al., 2023b). In this paper, we refer to Zhong et al. (Zhong et al., 2023) and use the codebase from Li et al. (Li et al., 2025), as it is an accelerated version.
@@ 4.1.5. Implementation Details
[…]
@@
[…]
@@
[…]
TREC DL 19 nDCG@10==0.650 Random Noise 0.735 0.023 0.046 0.037 1.000 0.984 3469.2 0.1
@@ 4.3.2. Black-box Attack Performance
It is worth noting that, in the literature (Wu et al., 2023; Liu et al., 2023b, 2024, a), black-box attacks typically involve distilling a surrogate model first, then using white-box methods to attack that surrogate model. However, distilling LLMs like LLaMA is too computationally intensive in our experiments. Therefore, to ensure fairness across all target models and to reduce the computational workload, we refrain from training surrogate models. Instead, we follow the approach in  (Li et al., 2025) and directly use the adversarial documents generated by various attack methods during attacks on SimLM and apply them to the target models.